OpenStudy (anonymous):

please help what is in the middle of 1/3 and 2/3

OpenStudy (turingtest):

1/3=2/6 2/3=4/6 3/6=1/2 is in the middle of these

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

use the linear graphics

OpenStudy (anonymous):

k